Variable costs per unit of production in 2019 amounted to 200 rubles per unit.

In total, 10000 items were manufactured.

Fixed costs for the year amounted to 20,000 rubles.

It is supposed to establish a 40% mark-up on products after manufacturing. Calculate your unit cost.

The selling price of the product.

The company's revenue from the sale of these 10000 products.

Calculate the company's gross profit, profit after tax, and gross margin (profit margin)

Tax -20 %


1
Expert's answer
2021-01-19T07:42:51-0500

Variable cost per unit = 200 rubles

Total units manufactured = 10,000 units

Total variable cost (TVC) = 200 * 10,000 = 2000,000 rubles

Fixed cost (FC) = 20,000 rubles

Total cost (TC) = TVC + FC = 2000,000 + 20,000

TC = 2,020,000 rubles

Company's revenue = 140% * 2,020,000 = 2,828,000 rubles

Selling price per unit = 2,828,000/10,000

= 282.8 rubles

Gross profit = TR - TC = 2828000-2020000 = 808,000

Tax = 20% * 808,000 = 161,600

Profit after tax = 808,000-161600 = 646400

Gross margin = 808000/2828000 *100% = 28.57%